### v1

# Relation of Continuous Chirality Measure to Spin and Orbital Polarization, and Chiroptical Properties in Solids

Abstract
Chirality introduces intriguing topological, electronic, and optical properties to molecules and solids. In this work, we investigate the influence of structural chirality on spin and orbital polarization as well as optical activity through first-principles calculations and continuous chirality measure (CCM). By using chiral selenium (Se) and 2D hybrid perovskites as examples, we demonstrate that chirality continuously modifies spin-orbit splitting and orbital angular momentum (OAM) polarization. We establish a direct relation between chirality transfer across organic-inorganic interfaces and inversion symmetry breaking, which induces Rashba-Dresselhaus spin splitting in hybrid perovskites. Additionally, we examine the effects of chirality on circular dichroism (CD) and the circular photogalvanic effect (CPGE), demonstrating how the continuous tuning of chirality dictates their magnitude and anisotropy. Our findings highlight that applying hydrostatic pressure can effectively tune the CCM, thereby enhancing the chirality-induced effects such as spin selectivity. By integrating CCM with electronic structure calculations, we present a predictive strategy for designing chiral materials with tailored optical and spintronic functionalities.
I Introduction
Structural chirality lacks spatial inversion and mirror symmetries, resulting in distinct electronic and optical responses [1]. This inherent asymmetry plays a crucial role in optical orientation, where chiral materials interact differently with left- and right-circularly polarized light. Such interactions give rise to circular dichroism (CD) [2] and circularly polarized luminescence (CPL) [3, 4]. The ability to control light-matter interactions through chirality enhances spin-dependent optical functionalities in modern optoelectronic applications. Besides optical activity, chirality has important implications in quantum transport; for example large spin polarization can be generated by running charges in nonmagnetic low-spin-orbit chiral materials, known as the chirality-induced spin selectivity (CISS) effect [5, 6]. As such effect implies spontaneous time-reversal symmetry broken in absence of permanent magnetization, the mechanism behind the CISS effect is still debated. Since its discovery in 2012 [5], researchers have proposed different theories, but no clear agreement has been reached [7, 8, 9]. Additionally, chirality plays a key role in orbitronics, an emerging field that utilizes electron orbital angular momentum (OAM) [10, 11, 12, 13] as an information carrier. The broken inversion symmetry in chiral materials enables coupling between OAM and spin angular momentum (SAM) [14, 15, 16, 17], leading to chiral orbital textures and potentially large orbital Hall conductivity (OHC) [13, 18]. By integrating these effects, chiral materials serve as a versatile platform for next-generation optoelectronic, spintronic, and orbitronic applications, enabling advances in quantum information science and low-power electronics [19, 20].
Quantifying chirality is critical to understanding its effects on electronic and optical properties. The Continuous Chirality Measure (CCM) offers a reliable way to do so. Unlike binary classifications of chirality, CCM provides a continuous scale to evaluate the degree of chirality in a given structure. For example, Hausdorff distances and CCM have been applied to crystal structures to distinguish chiral space groups and analyze their influence on band structures and Berry curvatures [21]. In another approach, graph-theoretical chirality (GTC) [22], based on torsion in differential geometry, has been proposed to describe chirality across different length scales, offering direct correlations with optical activity and chiral meta-materials. Beyond structural characterization, CCM has emerged as a crucial parameter in describing the electronic properties of chiral materials, in particular to their spin and orbital textures, as we will show in this work. In systems with the presence of CISS, a larger CCM often correlates with stronger spin-polarization (or spin-splitting in k space), suggesting a direct impact on spin-selective transport. Similarly, in chiral crystals, OAM textures emerge, and as we will show later, their properties can be quantitatively related to CCM. Thus, CCM not only measures chirality geometrically but also predicts emergent quantum properties.
CCM plays a crucial role in linking structural chirality to chiroptical properties such as CD and the Circular Photogalvanic Effect (CPGE). Since CD reflects differential absorption of left- and right-circularly polarized light, higher CCM intuitively correlates with a stronger CD signal. This is because a higher CCM may correspond to more pronounced chiral electronic states or electronic transition properties, thereby enhancing the optical response. Recent studies on chiral perovskites demonstrate that tuning molecular-level chirality through structural engineering enhances CD, reinforcing the connection between CCM and chiroptical properties [23]. Although CD is widely used to characterize chirality in molecules and materials, CPGE has been used to probe the topological charge of Weyl points in chiral materials [24, 25], therefore its precise relationship with CCM needs further experimental and theoretical investigation. On the other hand, CPGE offers a unique probe of chirality-sensitive charge dynamics, as its response is inherently linked to broken crystal symmetry and spin-orbit coupling (SOC) effects such as Rashba and Dresselhaus interactions [26, 27]. Since chirality breaks inversion symmetry, these SOC-induced band splittings can lead to spin-polarized photocurrents under circularly polarized light. This makes CPGE a promising tool for characterizing chiral materials, providing complementary insights beyond traditional chiroptical measurements like CD, particularly in chiral semiconductors and hybrid 2D-Perovskites [27].
The remainder of this paper is organized as follows. In Section II.A, we introduce the definition of the CCM for solid-state crystals and two-dimensional hybrid perovskites, establishing a quantitative framework for measuring the degree of chirality and exploring its correlation with spin and orbital textures in these systems. In Section II.B, we investigate the influence of chirality on optical responses, specifically CD and absorption dissymmetry factor ($g_{CD}$). We discuss how symmetry considerations, including inversion and mirror symmetry, govern these optical effects and further examine the role of strain in modulating CD. In Section II.C, we analyze the relationship between chirality and CPGE, highlighting the impact of SOC and Rashba-Dresselhaus effects on the observed trends. In Section II.D, we explore the transfer of chirality across organic-inorganic interfaces in hybrid perovskites, emphasizing how this process modifies electronic and optical properties, particularly in enhancing CD and CPGE. Finally, in Section III, we summarize our findings and propose future directions for leveraging chirality in the design of next-generation optoelectronic and spintronic materials.
II Result and Discussion
II.1 Continuous Chirality Measure
The CCM has been commonly used to quantify chirality in molecules and was recently adapted to describe periodic crystals [21]. CCM is defined by the mean-squared distance between atomic positions in the chiral structure and their closest equivalent positions in a symmetric reference structure. This approach ensures that the measure effectively captures the degree of chirality. The unnormalized chirality measure is defined as:
$\tilde{S}^{2}(G)=\sum_{i=1}^{n}\left\|p_{i}-p_{i}^{\text{sym}}\right\|^{2},$ (1)
where $p_{i}$ are the atomic positions in the target chiral structure, and $p_{i}^{\text{sym}}$ are the nearest equivalent positions in an achiral reference space group such that $\tilde{S}^{2}(G)$ is minimized.
To compare different structures, the measure is normalized as:
$S^{2}(G)=\frac{\tilde{S}^{2}(G)}{N}\times 100,$ (2)
where $N$ is a normalization factor accounting for the difference between the achiral and chiral structure:
$N=\sum_{i=1}^{n}\left\|p_{i}^{\text{chiral}}-p_{i}^{\text{sym}}\right\|^{2}.$ (3)
The positions of the chiral reference structure ($p^{\text{chiral}}$) are determined such that $N$ is maximized. The normalization allows $S^{2}(G)$ to vary from 0 (no chirality) to 100% (fully chiral). An ideal case system to study the change in CCM is trigonal selenium (Se).
[…]
Figure 1: Spin and orbital texture of chiral Se as a function of chirality. (a) Chiral Se Structure with $P3_{1}21$ symmetry. (b) Normalized Continuous Chirality Measure ($S^{2}$) as a function of the $3a$ Wycoff positions for $P3_{1}21$ symmetry. (c) and (e) are the Se band structure with bands colored by $\langle S_{z}\rangle$ and $\langle L_{z}\rangle$ respectively. (d) Fermi-Dirac average of $\langle|L_{z}|\rangle$ and spin splitting C as a function of chirality. (f) Se Spin texture centered at the H high-symmetry point in the $k_{y}$-$k_{z}$ plane parallel to the chiral axis z.
A second prototypical chiral system we investigate is two-dimensional (2D) hybrid organic-inorganic perovskites (HOIPs). Incorporation of chiral organic cations into 2D HOIPs results in a transfer of structral chirality to inorganic lattice, by inducing chiral distortion [28]. For example, R-(+)- and S-(-)-1-(1-naphthyl)ethylammonium (R/S-NEA) cations have been incorporated into layered lead bromide perovskites, forming (R/S-NEA)2PbBr4 structures (R/S-NPB). These chiral spacers induce symmetry-breaking distortions in the inorganic framework through asymmetric hydrogen bonding, an effect not observed in the racemic system. This results in the P21 chiral symmetry, leading to CD and Rashba-Dresselhaus spin splitting. To investigate the role of chirality transfer we generate configurations through a structural interpolation between the centrosymmetric structure and a fully chiral-distorted state. To simplify structural interpolation, we replace the chiral molecules with Cs+, which also removes any electronic effects resulting from the molecules (we note no Cs related states close to band edges). The initial reference structure (CCM = 0%) is an undistorted perovskite lattice and corresponds to the inorganic sublattice of racemic-NPB. The final structure (CCM = 100%) is the experimentally observed distorted state, where chiral organic cations cause a clear helical deformation in the inorganic layers, breaking mirror and inversion symmetry. To distinguish the hybrid structure with molecules and the interpolated structure with Cs, we adopt the notation of 2D-NPB and 2D-Cs-PB, respectively. To obtain structures with intermediate degree of chirality, we use a simple linear interpolation method based on atomic positions and lattice parameters. This method gradually changes the reference structure into the chiral distorted structure by smoothly adjusting atomic positions, bond angles, and torsional distortions along a defined path. Specifically, the interpolation follows:
$R_{j}(\lambda)=R_{j}^{(i)}+\lambda(R_{j}^{(f)}-R_{j}^{(i)})$ (4)
where $R_{i}(\lambda)$ is the the atomic position of the $j$-th atom at an interpolation step (ranging from 0 to 1), $R_{j}^{(i)}$ is the atomic position in the initial centrosymmetric reference structure, and $R_{j}^{(f)}$ is the final experimental chiral structure.
II.2 Spin and Orbital Texture
One key property of interest in chiral structures is the presence of orbital and spin polarization and textures, thus we systematically investigate how these properties change with CCM. We find both orbital and spin textures are strongly correlated with the degree of chirality. The Se crystal structure with SG P3121 possesses 3-fold screw rotation {$C_{3z}$mid 0, 0, $\frac{1}{3}$}, and 2-fold rotations {$C_{2y}$mid 0, 0, $\frac{1}{3}$} and {$C_{2,110}$mid 0, 0, 0} as the symmetry operators. These symmetries impose constraints on the SAM and OAM in momentum space, causing certain directional components to vanish along specific high-symmetry paths. Specifically, along the $\Gamma$–A and K–H lines, $\mathrm{S}_{x,y}(\mathbf{k})$ and $\mathrm{L}_{x,y}(\mathbf{k})$ vanish due to symmetry constraints, while $\mathrm{S}_{z}(\mathbf{k})$ and $\mathrm{L}_{z}(\mathbf{k})$ are nonzero with z chosen to be along the chiral axis. In contrast, along the $\Gamma$–K–M and A–H–L lines, $\mathrm{S}_{x,y}(\mathbf{k})$ and $\mathrm{L}_{x,y}(\mathbf{k})$ are nonzero, whereas $\mathrm{S}_{z}(\mathbf{k})$ and $\mathrm{L}_{z}(\mathbf{k})$ vanish (as shown in Figure 1 and SI Figure S1). We mainly focus on the SAM and OAM distribution at the conduction band minimum (CBM) of the H point, as it is a crucial Weyl point [14]. The SAM and OAM textures along the high-symmetry lines passing through the H point are shown in the Figure 1c and e. The results are consistent with the analysis based on symmetry constraints. The complete SAM and OAM-resolved band structures and textures of Se can be found in SI Figures S1 and S4. In general, we found a good correlation between spin and orbital textures in k-space in chiral systems.
Correlation of spin and orbital textures for Se and 2D-NPB ar highlighted in Figure 2. A persistent spin helix (PSH) like spin texture can be seen for both the spin and orbital textures in 2D-NPB, Figure 2a and c. Interestingly, the orbital texture flips sign around below 0.12 Å-1, becoming anti-parallel to the spin texture. Se has a more radial spin texture in the plane normal to the screw axis as discussed below and we see the similar pattern in the orbital texture, Figure 2b and d. In Se, perpendicular to the screw axis there is no radial spin texture and a more PSH like spin texture exists perpendicular to the plane, Figure 1f. Normal to the screw axis, there is no out of plane component to the spin or orbital textures for both Se and the 2D-Cs-PB system’s orbital and spin textures, indicated by the absence of any color in Figure 2.
Figure 2: Spin and orbital texture of chiral solids. (a) Spin texture in the $k_{x}$-$k_{z}$ plane centered at $\Gamma$ high-symmetry point of the CBM in 2D-NPB (z is the layer-spacing direction) and (b) spin texture in the $k_{x}$-$k_{y}$ plane centered at H high-symmetry point of the CBM in Se, normal to the chiral axis z. Orbital texture in the same planes for (c) 2D-NPB and (d) Se. Color bars indicate the out-of-plane component of spin/orbital component.
Figure 3: Spin texture analysis for Weyl, Rashba, and Dresselhaus contributions as a function of chirality for two systems. (a) 2D-NPB with major and minor $P2_{1}$ screw axis highlighted in green and red, respectively. Normalized magnitude of fitting matrix $C$ for (b) 2D-Cs-PB and (c) Se. Contributions from Weyl, Rashba, and Dresselhaus spin types to the spin splitting (d) 2D-Cs-PB and (e) Se.
In order to quantify spin texture we employ an effective Zeeman Hamiltonian as in [29]:
$H^{\text{in}}=\frac{2g_{0}\mu_{B}}{\hbar}\boldsymbol{B}^{\text{in}}(\textbf{k})\cdot\boldsymbol{S_{\textbf{k}}}$ (5)
$\boldsymbol{B}^{\text{in}}(\textbf{k})=C\cdot\boldsymbol{\textbf{k}}$ (6)
$C=C_{s}+C_{v}+C_{t}$ (7)
where $\mu_{B}$ is the Bohr magneton, $g_{0}$ is the free electron $g$-factor, $S_{\textbf{k}}$ is the spin angular momentum, $\boldsymbol{B}^{\text{in}}(\textbf{k})$ is the internal magnetic field. Given the internal magnetic field, the coefficient tensor $C$ can be extracted through least-squares fitting. The $C$ tensor can then be decomposed, enabling the quantification of different contributions to the spin texture. The Weyl contribution ($C_{s}$) is the the trace, the Rashba contributions ($C_{v}$) is the traceless antisymmetric part, and the Dresselhaus contributions ($C_{t}$) is the symmetric part [29]. For both systems we tested in Figure 3b and c, the magnitude of $C$ (the magnitude of k-dependent internal magnetic field $\boldsymbol{B}^{\text{in}}$) increases with chirality. This directly reflects the relation between symmetry broken and strength of spin-orbit field. While overall the magnitude of $C$ increases, the relative contributions vary significantly between the 2D-Cs-PB and chiral Se.
In chiral Se, the spin texture was found to consist of primarily the Weyl and Dresselhaus terms as shown in Figure 3e. At low chirality, the Weyl contribution is dominant, and the Dresselhaus component increases with increasing chirality. The trend is consistent with a qualitative analysis of the spin texture normal to the chiral axis (see Figure 2b, which shows a predominantly radial pattern with a certain degree of radial-tangential mixing). According to the classification summarized by Zunger et al. [30], such a texture indicates a hybrid spin polarization character that origins from a mixture of Weyl and Dresselhaus terms. However, the increasing Dresselhaus contribution may be missed from a qualitative analysis. Similar to the spin texture, the orbital texture also exhibits the radial-tangential mixing for all chirality degree, see SI Figure S4.
In the 2D-Cs-PB structure at low chirality, all three spin texture types are present in comparable magnitudes (Figure 3d). Similar to Se, the Weyl contribution decreases as chirality increases. However, in contrast to Se, the Rashba contribution increases when chirality increases. Due to the large Rashba and Dresselhause components, the resulting spin texture resembles persistent spin helix (PSH) [29] . At the highest chirality there is an equal contribution of Rashba and Dresselhaus resulting in a nearly perfect PSH spin texture as shown in Figure 3d. Similar to Se, the orbital texture closely resembles the spin texture as shown in SI Figure S3. Since, the degree of chirality in the halide perovskite is dictated by the interaction with chiral molecules, one can experimentally tune spin and orbital textures by engineering chiral molecules.
II.3 Effect of Chirality on CD
Optical activity such as Circular Dichroism (CD) is a direct probe for chirality. We will discuss the quantitative relation between CD and chirality degree in the following. The CD response was computed based on the Berry phase formalism as detailed in the previous study [31], including both magnetic dipole (MD) and electric quadrupole (EQ) contributions. The latter is particularly important for anisotropic solids [32]. We note that excitonic effects are not included in the current study, where we mainly focus on the above-bandgap transitions. We have shown previously the CD responses of above-bandgap transition in 2D hybrid perovskites are well reproduced with the current implementation [31, 32]. In the following, first, we confirm the magnitude of CD generally increases with increasing chirality degree in both systems. We next will focus on the chirality dependence in various aspects of CD, including anisotropicity, MD/EQ contribution, spin/orbital components, and $g_{CD}$.
Figure 4: CD of the 2D-Cs-PB system as a function of chirality. (a) CD response in the x-direction and (b) yz direction. (c) $g_{CD}$ in the yz direction. CD from different contributions, (d) magnetic diploe (MD), (e) electric quadrupole (EQ) and (f) spin with one order lower than the other components. Note that a rigid shift of spectra has been applied to align the bandgap with the 100% chirality.
In the 2D-Cs-PB system, the $P2_{1}$ screw axis is in-plane along the y direction formed by the equatorial Br ions (Figure 3a), and leads to highly anisotropic CD responses as shown in the SI, Figure S7. The maximum response appears along the yz direction and increases with chirality degree monotonically, while the other directions have a more subtle relation to chirality. This can be rationalized since the yz direction not only includes the chiral distortions introduced into the equatorial Br but also includes chiral distortions to the axial Br atoms, see SI Figure S7. Interestingly, in all other directions minimal change in CD is observed until 100% of the chiral distortion is introduced. Additionally, we can remove the contributions from optical absorption to CD responses, by computing $g_{CD}$ which is defined as $g_{CD}=CD/A$ where $A$ is the optical absorbance. In the 2D-Cs-PB systems we found $g_{CD}$ has similar trends for CD and chirality, as shown in Figure 4c. This is due to a nearly constant absorption with respect to chirality.
To further examine the origin of CD increases with chirality, we decompose the contributions to CD into magnetic dipole, electric quadrupole, and spin contributions, as shown in Figure 4d, e, and f. Here we focus on the discussion on the yz direction since this direction has the largest CD response, although similar trends hold for other directions. The main contribution to CD is from the magnetic dipole (Figure 4d), with minor electric quadrupole contributions (Figure 4e). The spin contribution is found to be 1 order of magnitude lower than the orbital contributions (MD+EQ). Both the spin and MD contributions increase with increasing chirality, while no clear trend exists for the electric quadrupole.
Figure 5: CD and $g_{CD}$ of the bulk Se system as a function of chirality. CD of bulk Se with different chiralities, (a) perpendicular to and (c) along the screw axis in z. (b) $g_{CD}$ of Se with different chiralities perpendicular to and (d) along the screw axis. To enable clearer comparison, CD spectra were rigidly shifted along the energy and intensity axes for visualization purposes. The energy axes of all other chiralities were referenced to the spectrum of the configuration with $u$=0.218 ($S^{2}=48\%$).
In contrast to the 2D-Cs-PB system, Se only has diagonal contributions to the CD tensor due to the $D_{3}$ point group symmetry. The components perpendicular to the chiral axis (xx and yy) are equal due to symmetry constraints but differ parallel to the chiral axis (zz), reflecting the intrinsic anisotropy in the optical response. With increasing chirality, CD signals exhibit an increasing trend both perpendicular and parallel to the screw axis. Again we find the MD contribution being dominant, but in contrast to the 2D-Cs-PB, the MD contribution is opposite in sign to the EQ, SI Figure S6).
If we compare $g_{CD}$ perpendicular (xx) and parallel (zz) directions to the chiral axis in Figure 5b and d, we find that along x $g_{CD}$ increases significantly with increasing chirality, similar to 2D-Cs-PB, while along z (chiral axis) $g_{CD}$ remains relatively unchanged with chirality. At low energies close to the band edge, both CD and absorption in the zz component remain small and show little change, resulting in an almost constant $g_{CD}$. At higher energies, however, $g_{CD}$ still decreases as chirality reduces.
The key difference between Se and 2D-Cs-PB can be attributed to changes in the chemical bonds with chirality. In 2D-Cs-PB, the chirality is mainly introduced by the tilting of the corner-sharing $PbBr_{6}$ octahedra, and thus leads to significantly less change in the Pb-Br bonds. However, as the Wyckoff position parameter $u$ increases, the helical atomic chains of selenium gradually overlap within the cross-sectional (xy) plane, reducing the structural chirality. This process is accompanied by a transition in the interchain interaction—from weak van der Waals forces to stronger covalent bonding—as the atomic chains come into closer proximity. This is further reflected in the band structures (SI Figures S1 and S2), where 2D-Cs-PB exhibits negligible changes with varying chirality, unlike selenium which shows significant band evolution.
Additionally, we note that such continuous tuning of chirality in Se can be realized by applying strain [21, 33] as shown in SI Figure S8. As the pressure increases from 0 to 1 and then to 4 GPa, the corresponding chiralities decrease from 48%, to 35%, and finally to 22%, which is determined by the $3a$ Wycoff position parameter $u$ for $P3_{1}21$ symmetry. The variation of CD and $g_{CD}$ under pressure as a function of chirality in SI Figure S8 is similar to the unstressed case discussed earlier.
II.4 Effect of Chirality on CPGE
Figure 6: CPGE analysis as a function of chirality for 2D-Cs-PB. (a) CPGE tensor as a function of chirality. Contribution from (b) velocity difference and (c) dipole-matrix elements. Note in panels (a-c), a rigid shift has been applied to the excitation energies to align the bandgap with the 100% chiral structure. (d) Bandstructure of 2D-NPB with color indicating contributions from Pb, Br, and organic (Org.). Correlation of CPGE to (e) the percentage of Rashba SOC $C_{v}$ from $C$ tensor and (f) $\alpha$ coefficient fit using the Rashba Hamiltonian, linear fit plotted as a red dashed line.
Circularly-polarized light induced DC photocurrent is present in certain non-centrosymmetric materials, knonw as the circular photogalvanic effect, CPGE. Although chirality is not required for CPGE, structures with chirality induce CPGE since they lack both inversion and mirror symmetries. Experimentally, CPGE is often used to probe the amount of spin splitting, the type of spin-orbit coupling [26, 27], or non-trivial topology [24, 25]. We discuss the physical mechanism and quantitative relationship of CPGE and chirality as follows, the magnitude and symmetry of CPGE, the contributions from individual components, and its relation to SOC.
As demonstrated in Figure 3b and c, both systems exhibit the SOC field increase with increasing chirality, therefore a systematic CPGE change at the band edge may be expected. However, CPGE responses away from the band edge may have different origins, which don’t result from spin splitting alone. Specifically, based on the Kubo formula of CPGE with relaxation time approximation in Eq. 9, both light absorption (electronic transition dipoles $r^{k}_{\boldsymbol{k},nm}$) and electron-hole group velocity difference ($\Delta^{i}_{\boldsymbol{k},nm}$) determine the magnitude of CPGE. We focus on analyzing the 2D-Cs-PB system as an example in the following. Similar analysis and discussions for CPGE of bulk Se can be found in SI Figure S9.
First, CPGE in the 2D-Cs-PB system increases at all excitation energies as a function of chirality, as shown in Figure 6a. In this system, an increase in spin-splitting, electron-hole velocity difference, and dipole-transitions coexist, as seen in Figure 6. We observe the velocity difference only becomes substantial at higher excitation energies. In contrast, the dipole-transition is also significant near the band edge. The combined results show an increase in CPGE with chirality both near and above the band-edge.
In general, increasing chirality results in an increase in CPGE. The one exception is a slight reduction of CPGE between 75% to 100% chirailty at the maximum of the first peak, as shown in Figure 6a. There is also a reduction of electron-hole velocity difference and a peak shift to lower energies, as seen in Figure 6b. Additionally, the sign change in the dipole moment also moves towards lower energies. The combination of the velocity difference and the dipole transition explains the the slight reduction of the CPGE at the highest chirality.
In the 2D-Cs-PB system, the symmetry of the CPGE tensor is quite different from CD. The CD response is highly anisotropic along and perpendicular to the $P2_{1}$ screw axis, but there exists no such relation for CPGE. Instead, the CPGE parallel and perpendicular to the screw axis is roughly equal and opposite, as seen in SI Figure S10. Both directions have a strong correlation with chirality increasing with increasing chirality.
Additionally, we study the band edge CPGE response relation to the Rashba spin splitting in two ways. Specifically, we take the maximum of the first peak in Figure 6a and plot against the SOC properties in Figure 6e and f. There is strong correlation with both results described in Section IIB extracted for Rashba SOC $C_{v}$, and the $\alpha$ parameter of the Rashba Hamiltonian:
$E^{\pm}(\boldsymbol{k})=\frac{\hbar^{2}\boldsymbol{k}^{2}}{2m^{*}}\pm\alpha|\boldsymbol{k}|,$ (8)
where $m^{*}$ is the carrier effective mass and $E^{\pm}$ is the energies of the spin-split bands. For both methods the general trend is similar, namely CPGE increases with increasing Rashba splitting, although there is a slightly better correlation between Rashba SOC percentage and CPGE, Figure 6e.
II.5 Chirality Transfer through Electronic and Structural Effects on Optical Activity
The CD response in 2D chiral perovskites can have contributions from both distorted inorganic lattice (due to interactions with molecules) and chiral molecules themselves. To decompose the relative contributions we examine four systems: the molecule system alone (Molecule), the chiral perovskite sublattice (Inorganic), the combined molecule and inorganic sublattice (2D-NPB) and the combined system where we restore the original perovskite symmetry by removing chirality from the inorganic layer (NC-NPB). Looking at the separate contributions, the molecular component (Molecule) exhibits a CD response much larger than that of the inorganic component alone (Inorganic). However, when the molecular and inorganic parts are combined into a hybrid system, the overall CD signal is notably enhanced. The amplification of CD suggests a synergistic interplay between the structural chirality and the electronic contributions. This is further supported by the reduction in CD when the chiral distortions are removed from the inorganic sublattice (NC-NPB), Figure 6a. The enhancement of CD from the combination of structural distortion and electronic effects is consistent with previous work on remote chiarlity transfer from a chiral molecule to achiral perovskites [34]. Specifically, the electronic interactions, such as charge transfer or orbital overlap, along with the structural distortion introduced by molecule-inorganic lattice interactions, determine the overall CD response. These factors are critical to understanding the signatures of CD in the hybrid system.
We further analyze CD at 3.0 eV by projecting the pair of states in the transition matrix elements to atomic orbitals, as shown in Figure 6c. The analysis reveals three dominant transitions that correspond to the main peak at 3.0 eV. The strongest contribution comes from the transition from Br-$p$ to Pb-$p$ orbitals, followed in intensity by C-$p$ to Pb-$p$ and Br $p$ to C-$p$. Interestingly, the Br-$p$ to C-$p$ sign is opposite to the other main contributions and results in CD reduction. However, the most dominant contribution comes from the states of the inorganic lattice. Such an analysis highlights the enhancement of the CD response in inorganic lattices that require wavefunction overlaps with organic molecules.
Figure 7: CD and CPGE in 2D-NPB and their transition analysis. (a) CD and (b) CPGE in 2D-NPB for with separated contributions from the inorganic, molecule, and with the molecules present but no chirality in the inorganic layer (NC-NPB). Note all energies have been shifted to the minimum bandgap indicated by the dashed line. Transition analysis for at 3.0 eV for CD (c) and for CPGE at 2.4 eV (d) in 2D-NPB.
In contrast to CD, the CPGE is dominated mainly by contributions from the inorganic sublattice. As seen in Figure 7b, the inorganic contribution almost reproduces exactly the combined organic-inorganic response. This can be explained by the composition of the band-edge states which are mostly inorganic in character. Another noticeable difference between CD and CPGE is the onset of CPGE started at much lower excitation energies i.e. 1.5 eV compared to CD (its onset above 2 eV). Although there may be less absorption at the band edge, there is larger spin splitting giving larger velocity difference, which results in nonzero CPGE at lower energies than CD.
Similar to CD analysis, we project the transition matrix elements onto the specific transitions and find the majority of the CPGE is from the inorganic (Br-$p$ to Pb-$p$), with minor contributions from C-$p$ to Pb-$p$. More importantly, when removing the chiral distortion from the inorganic sublattice but keeping organic molecules present, we see almost no CPGE. This indicates that the chiral distortion to the inorganic sublattice is critical for controlling CPGE in 2D-hybrid perovskites.
III Conclusion
In this work we reveal the correlation between the continuous-chirality measure (CCM) and spin, orbital polarization, and optical activity of solids from first-principle calculations. First, through quantitative analysis of spin textures, we find that the spin orbit field or effective internal field monotonic increases with increasing chirality. The specific SOC components vary rather differently among the two systems. Meanwhile, we find that orbital angular momentum textures closely resemble those of spins. The optical activity such as circular dichorism is confirmed as a direct measure of chirality; however, we emphasize the anisotropy of CD needs careful consideration. Furthermore, through examination of different components of CD, we find MD contributions show correlation with chirality, but not EQ and spin. Furthermore, the absorption normalized CD, $g_{CD}$, shows increases with chirality in certain directions, but not all. In contrast, the CPGE does not show a consistent correlation with chirality. Even though spin splitting may be enlarged with chirality, other factors such as absorption or electron-hole velocity difference may reduce the CPGE response. In addition to the role of chirality, we investigate the contributions to CD and CPGE from different components of the 2D-HOIP. We demonstrate the significant role of synergistic relationship between structural chirality and electronic effect through organic-inorganic interactions in producing the maximum CD. In contrast, the chirality transfer to the inorganic sublattice is more dominant for CPGE, with little direct contribution from molecules. This work provides critical insights into the role of chirailty on spin, orbital textures, chiral-optical properties, which can be guidelines of controlling these properties through strain engineering or chirality transfer at interfaces.
IV Computational Methods
The electronic structure and geometry optimization calculations were performed in the open-source plane-wave DFT software JDFTx [35]. We employed norm-conserving pseudopotentials from PseudoDojo [36], using scalar-relativistic and fully relativistic treatments for calculations without and with spin-orbit coupling, respectively, with the plane wave basis set defined by a kinetic energy cutoff of 60 Ha for Se and 50 Ha for the 2D-perovskites. The exchange-correlation functional used in all calculations was the Perdew-Burke-Ernzerhof (PBE) generalized gradient approximation [37]. The initial structure of hybrid perovskite is from Ref. [28], and Se is from the Materials Project [38]. See Supplemental Material for more structural information and additional computational details. For Brillouin zone sampling in self-consistent calculations, we used a $\Gamma$-centered k-mesh. Specifically, a $5\times 5\times 1$ mesh was used for 2D hybrid perovskites, while a $12\times 12\times 12$ mesh was used for Se. Self-consistent calculations were followed by non-self-consistent calculations with high density k-points for the evaluation of matrix elements.
We employed random sampling within the first Brillouin zone. In a crystalline system, $L$, $Q_{\mu\nu}$ and $R$ elements often exhibit sharp peaks rather than varying smoothly with $k$, thus special treatment is required to ensure computational accuracy [31].
The CPGE tensor was implemented in an in-house code that is interfaced with JDFTx. CPGE in this work was computed from the following, mainly for light-induced injection current, derived from the second-order perturbation theory under the relaxation time approximation:
$\beta_{ij}=\frac{e^{3}\pi}{N_{\boldsymbol{k}}V\hbar}\epsilon_{jkl}\sum_{\boldsymbol{k},n,m}f^{\boldsymbol{k}}_{nm}\Delta^{i}_{\boldsymbol{k},nm}r^{k}_{\boldsymbol{k},nm}r^{l}_{\boldsymbol{k},mn}\delta(\hbar\omega-E_{\boldsymbol{k},nm})$ (9)
where $n$ and $m$ are band indices and $i,j,k$ and $l$ are Cartesian directions. Here, $f_{nm}$ is the occupation difference, $E_{\boldsymbol{k},nm}$ is the energy difference, $r^{k}_{\boldsymbol{k},nm}$ are the dipole-matrix elements, and $\Delta^{i}_{\boldsymbol{k},nm}$ are the difference in velocity-matrix elements of band $n$ and $m$ at k-point $\boldsymbol{k}$. All these elements are computed from DFT directly, with Monte-Carlo k-point sampling, without Wannier interpolation. We then benchmark against previous results where Wannier interpolation was used, see SI Figure S11. The BZ sampling technique similar to the CD calculations was implemented for CPGE.
acknowledgement
We acknowledge support for the theoretical and code development of optical activity, CPGE, and CCM by the computational chemical science program within the Office of Science at DOE under grant No. DE-SC0023301. We acknowledge the support for materials theory application as part of the Center for Hybrid Organic-Inorganic Semiconductors for Energy (CHOISE), an Energy Frontier Research Center funded by the Office of Basic Energy Sciences, Office of Science within the US Department of Energy (DOE). This research used resources of the Scientific Data and Computing center, a component of the Computational Science Initiative, at Brookhaven National Laboratory under Contract No. DE-SC0012704, the National Energy Research Scientific Computing Center (NERSC) a U.S. Department of Energy Office of Science User Facility operated under Contract No. DE-AC02-05CH11231. This work used the TACC Stampede3 system at the University of Texas at Austin through allocation PHY240212 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services and Support (ACCESS) program [39], which is supported by US National Science Foundation grants No. 2138259, No. 2138286, No. 2138307, No. 2137603, and No. 2138296.